# A Novel Sampled Clustering Algorithm for Rice Phenotypic Data

### v1

@@ 5.2 Ideal Number of Clusters
Figure 1: First 30 eigenvalues obtained using natural exponential function and base “$30$” exponential function for building the similarity matrix.
Determining the ideal number of clusters is always a dataset specific task, and we have many methods to approximate this number. A commonly used method is the Eigen-Gap heuristic (Von Luxburg (2007)) tailored for Spectral Clustering. This heuristic says that given $\lambda_{i}$’s as the eigenvalues of the Laplacian matrix, $k$ is the ideal number of clusters when the eigenvalues $\lambda_{1}$, $\lambda_{2}$, .., $\lambda_{k}$ have small magnitude while $\lambda_{k+1}$ has comparatively much higher magnitude. Figure 1 displays the $30$ smallest eigenvalues derived from the Laplacian matrix obtained during the application of our locally scaled base “$30$” exponential Spectral Clustering on the rice data. From this figure, we see a jump at 1, 2, 3, 5, and 10. To double-check our result, we also apply the Spectral Global Silhouette method proposed in Mur et al. (2016). This method uses the concept of local scaling, as discussed earlier, and gives 5 as the ideal number of clusters. Plant biologists are typically interested in more number of clusters, and hence, we choose 5, 10, 15, and 20 as the ideal number of clusters for our experimentation.
6 Analysis
In this section, first, we demonstrate that eigenvalues of the Laplacian in (3) reduce when we use an exponential function with base $a$ (i.e., (7)), instead of a natural exponential function (i.e., (2)) in building the similarity matrix. Second, we check the quality of our samples via a sampling analysis.
6.1 Eigenvalue Analysis
Table 2 lists the first 30 smallest eigenvalues of the Laplacian matrix obtained from both the natural exponential function similarity matrix and base “$30$” exponential function similarity matrix. Additionally, Figure 1 plots these eigenvalues. From both the table and the figure, we observe that the eigenvalues associated with the latter similarity matrix are closer to zero as compared to the eigenvalues associated with the former similarity matrix. This finding validates our Conjecture 3.3. Thus, using Cheeger’s Inequality (8), we infer that our novel approach of constructing the similarity matrix using a base “$30$” exponential function should lead to improved clustering. This is supported experimentally in the next section.
Sr. No. 1 2 3 4 5 6 7 8 9
Base (e) Fun 0.0000 0.6653 0.8397 0.8677 0.9038 0.9159 0.9235 0.9402 0.9472
Base (30) Fun 0.0000 0.4272 0.6642 0.7078 0.7134 0.8025 0.8151 0.8261 0.8441
Sr. No. 10 11 25 26 27 28 29 30
Base (e) Fun 0.9538 0.9606 0.9916 0.9918 0.9926 0.9929 0.9934 0.9936
Base (30) Fun 0.8468 0.8761 0.9420 0.9439 0.9457 0.9468 0.9501 0.9506
Table 2: Small eigenvalues of Laplacian matrices obtained using different similarity matrices.
6.2 Sampling Analysis
In order to assess the quality of the obtained samples, we employ the estimation of the following metrics: the Population Total (PT) estimator and the Horvitz-Thompson (HT) estimator. We define the PT estimator for a specific characteristic as the summation of all the values associated with that characteristic. That is, for a characteristic $j$, it is computed as
$(Y_{PT})_{j}=\sum_{i=1}^{U}(\chi_{j}){}_{i},$ (17)
where U is the set of all species (or #(U)=n) and $(\chi_{j}){}_{i}$ is the value of the $j^{t}{}^{h}$ characteristic for the $i^{t}{}^{h}$ specie.
The HT estimator is defined as follows:
$(Y_{HT})_{j}=\sum_{i=1}^{S}\frac{(\chi_{j}){}_{i}}{\pi_{i}},$ (18)
where $S$ is the set of samples (or #(S)=N), $(\chi_{j}){}_{i}$ is defined as above, and $\pi_{i}$ denotes inclusion probability of $i^{t}{}^{h}$ specie. The sampling is better when the values of the HT estimators are closer to the PT estimators.
Characteristics PT HT (Max) HT (Median)
estimators estimators estimators
  1 9.037800e+03 9.099514e+03 9.032072e+03
  2 2.121750e+05 2.113126e+05 2.103986e+05
  3 3.057500e+04 3.120878e+04 3.089743e+04
  4 1.614470e+04 1.619068e+04 1.612586e+04
  5 5.661400e+03 5.625295e+03 5.652058e+03
  6 4.630100e+03 4.599130e+03 4.538882e+03
  7 1.865000e+05 1.885044e+05 1.852332e+05
  8 3.343930e+04 3.508258e+04 3.456751e+04
  9 1.024092e+05 1.034117e+05 1.026595e+05
  10 2.569800e+03 2.552394e+03 2.562038e+03
  11 4.660790e+04 4.639347e+04 4.629499e+04
  12 7.214200e+04 7.184603e+04 7.227474e+04
Table 3: PT and HT Estimators.
Table 3 lists the PT estimators, HT estimators when using the maximum function in (14), and HT estimators when using the median function in (14) for all the 12 characteristics. The first thing we observe from the table is that the PT estimators are close to both the sets of HT estimators, and hence, both the samplings are good. Second, we also observe that the PT estimators are closer to the HT estimators when using the median function than the HT estimators when using the maximum function.
To better demonstrate the second observation, we plot the difference between the PT and the HT estimators (when using both the maximum function and the median function) for all characteristics in Figure 2. We can notice from this figure that the median function leads to a smaller difference in 9 of the 12 characteristics as compared to the maximum function. Thus, median based sampling is better.
Figure 2: Difference between PT and HT estimator when sampled with maximum function and median function.
Old New
Clusters Distance Old SC New SC HC Vs Vs
HC % HC %
Euclidean 0.1249 0.1258 0.07
5 SqEuclidean 0.215 0.2242 0.1349 28.88 34.15
Correlation 0.22 0.229 0.1707
Euclidean 0.0952 0.1113 0.0662
10 SqEuclidean 0.1592 0.2002 0.1079 35.64 55.43
Correlation 0.1747 0.1987 0.1288
Euclidean 0.0841 0.0946 0.026
15 SqEuclidean 0.1342 0.1725 0.0677 51.55 72.33
Correlation 0.1517 0.1719 0.1001
Euclidean 0.0784 0.0874 0.0128
20 SqEuclidean 0.1109 0.1587 0.0432 83.35 103.40
Correlation 0.1454 0.1613 0.0793
Average percentage gain 49.86 66.33
Table 4: Silhouette values of clustering algorithms.
7 Results
In this section, we perform two kinds of experiments. In Section 7.1, we compare the different clustering algorithms without sampling. In Section 7.2, we compare the sampled clustering algorithms.
7.1 Clustering without Sampling
@@
Here, we compare three clusterings. First is the standard Spectral Clustering as described in Section 3.1 (also natural exponential Spectral Clustering), and used in Shastri et al. (2021). We refer to this as Old SC. Second is the locally scaled base “$30$” exponential Spectral Clustering as elaborated in Section 3.2. We call this the New SC. Finally, the third is Hierarchical Clustering, which is mentioned in Section 2. We refer to this as HC.
@@
The results of this comparison are given in Table 4. Here, the first column denotes the number of the clusters that are chosen based upon the previous analysis. The second column contains the distance metrics used to build the similarity matrix in the spectral clustering algorithms. Columns three through five list the silhouette values of the respective algorithms. Best values in a cell are highlighted in bold. Finally, columns six and seven give the percentage gain of Old SC and New SC over HC, respectively. The best values are used to compute this gain. As evident, the New SC works the best here.
Old New Sampled Old New
Clusters Distance Sampled SC Sampled SC HC Vs Vs
HC % HC %
Euclidean 0.1218 0.127 0.1086
5 SqEuclidean 0.2111 0.2233 0.1426 25.61 30.08
Correlation 0.2276 0.2357 0.1812
Euclidean 0.0851 0.1043 0.0664
10 SqEuclidean 0.1602 0.1943 0.1 26.20 45.68
Correlation 0.171 0.1974 0.1355
Euclidean 0.0815 0.0858 0.0591
15 SqEuclidean 0.1348 0.1586 0.0917 11.01 15.10
Correlation 0.1573 0.1631 0.1417
Euclidean 0.0751 0.0774 0.0453
20 SqEuclidean 0.1115 0.1367 0.0856 6.68 14.81
Correlation 0.1405 0.1512 0.1317
Average percentage gain 17.37 26.42
Table 5: Silhouette values of clustering algorithms with sampling.
Clusters Distance New HC New Sampled
Sampled SC SC vs HC %
Euclidean 0.127 0.07
5 SqEuclidean 0.2233 0.1349 38.08
Correlation 0.2357 0.1707
Euclidean 0.1043 0.0662
10 SqEuclidean 0.1943 0.1079 53.26
Correlation 0.1974 0.1288
Euclidean 0.0858 0.026
15 SqEuclidean 0.1586 0.0677 62.93
Correlation 0.1631 0.1001
Euclidean 0.0774 0.0128
20 SqEuclidean 0.1367 0.0432 90.67
Correlation 0.1512 0.0793
Average percentage gain 61.24
Table 6: Silhouette values of New Sampled SC and HC.
7.2 Sampled Clustering
Here, we again compare the three clusterings (as in the previous section) but now after sampling, First, is the standard Spectral Clustering from Section 3.1 (natural exponential Spectral Clustering) along with the maximum function based Pivotal Sampling from Section 4, as used in Shastri et al. (2021). We refer to this as Old Sampled SC. Second is the locally scaled base “$30$” exponential Spectral Clustering from Section 3.2 along with the median function based Pivotal Sampling from Section 4. We call this New Sampled SC. Finally, as mentioned earlier, the third is the Hierarchical Clustering along with the median based Pivotal Sampling. We refer to this as Sampled HC. The results for this are given in Table 5. All columns here are the same as in Table 4. As evident, the New Sampled SC works best again.
To conclude, in Table 6, we compare New Sampled SC with HC. Thus, our new algorithm is substantially better than the current best and also computational cheaper.
8 Conclusions and Future Work
In one of our recent works, we had developed a probabilistically sampled (using pivotal sampling) and spectrally clustered algorithm for grouping soybean species. In this work, we improve this algorithm and apply it to rice species. The first two improvements involve use of better functions to build the similarity matrix in the spectral clustering algorithm. These design choices are supported via an eigenvalue analysis. The third improvement is using a more intuitive function (median) instead of a maximum function in computing the probabilities in the pivotal sampling algorithm. This design choice is supported via the computation of statistical measures.
[…]
@@ 8 Conclusions and Future Work
There are multiple future work directions here. First, in one of the seminal works (Ng et al. (2001)), the authors have listed sufficiency conditions for SC to work well. It would be very useful to translate those conditions to plant data. Second, it would be useful to experiment with other accurate clusterings and samplings (e.g., see Jain et al. (2021)). Third, although phenotypic characteristics are useful for clustering, genetic data of plant species carries more information. In our earlier work (Shastri et al. (2019)), we had explored the possibility of using genetic data for clustering, however, reduced data was used there. It would be interesting to experiment with the full data exhaustively.

### v3

@@ IV-B Normalization
Figure 2: Soybean: First 30 eigenvalues obtained using natural exponential function, base “$30$” exponential function, and base “$30$” locally scaled exponential function for building the similarity matrix.
IV-C Eigenvalue Analysis
Fig. 2 and 3 plot the eigenvalues for soybean and rice, respectively. These are first $30$ smallest eigenvalues of the Laplacian matrix obtained from similarity matrix built using natural exponential function, base “$30$” exponential function, and base “$30$” locally scaled exponential function.
These figures validate our Conjecture III.3. That is, the eigenvalues associated with base “$30$” exponential function are closer to zero as compared to the eigenvalues associated with the natural exponential function. Thus, as mentioned earlier, using Cheegers’s Inequality (8), base “$30$” exponential function should result in better clustering than the natural exponential function based clustering. This turns to be true experimentally, which we demonstrate in the results section.
Second, we further observe that, as claimed in Section III-B, eigenvalues corresponding to base “$30$” locally scaled exponential function are more closer zero than the prior two function choices. Thus, again by using Cheegers’s Inequality (8), this function should give the best clustering. This turns to be true experimentally as well, which we again demonstrate in the results section.
Figure 3: Rice: First 30 eigenvalues obtained using natural exponential function, base “$30$” exponential function, and base “$30$” locally scaled exponential function for building the similarity matrix.
V Results
As discussed in the introduction, we perform experiments on $2376$ soybean and $1865$ rice species. Determining the ideal number of clusters remains an open problem in SC. Based on inputs from plant biologists (which is based on the available number of species of each type) we cluster the soybean data into $10$, $20$, and $30$ groups, and the rice data into $5$, $10$, $15$, and $20$ groups.
To evaluate the quality of clustering we follow the standard definition of Silhoutte Value. This value for data point $p_{i}$ is given as
$s(p_{i})=\frac{b(p_{i})-a(p_{i})}{\max(a(p_{i}),b(p_{i}))},$ (13)
where $a(p_{i})$ denotes the average distance of $p_{i}$ to the points in its own cluster, while $b(p_{i})$ is the average distance of $p_{i}$ to points in its closest cluster.
Old Base New Old SC Base “$30$ New SC
Clusters Distance SC $30$” SC SC HC Vs SC Vs Vs
HC % HC % HC %
Euclidean 0.2422 0.2520 0.2562 0.2173
10 SqEuclidean 0.3836 0.3905 0.4066 0.3257 17.78 23.42 33.12
Correlation 0.3426 0.4020 0.4336 0.2307
Euclidean 0.2069 0.2148 0.0946 0.1833
20 SqEuclidean 0.2612 0.3191 0.3151 0.2095 24.69 58.8 63.67
Correlation 0.2313 0.3327 0.3429 0.0598
Euclidean 0.1783 0.1850 0.1815 0.1158
30 SqEuclidean 0.1538 0.2588 0.2865 0.1086 53.97 136.01 147.41
Correlation 0.1556 0.2734 0.2824 0.0535
Average percentage gain 32.15 72.74 81.40
TABLE III: Silhouette values of different clustering algorithms for Soybean.
Old Base New Old SC Base “$30$ New SC
Clusters Distance SC $30$” SC SC HC Vs SC Vs Vs
HC % HC % HC %
Euclidean 0.1249 0.1235 0.1258 0.07
5 SqEuclidean 0.215 0.2214 0.2242 0.1349 28.88 33.8 34.15
Correlation 0.2200 0.2284 0.2290 0.1707
Euclidean 0.0952 0.1103 0.1113 0.0662
10 SqEuclidean 0.1592 0.1952 0.2002 0.1079 35.64 51.55 55.43
Correlation 0.1747 0.1913 0.1987 0.1288
Euclidean 0.0841 0.0981 0.0946 0.026
15 SqEuclidean 0.1342 0.1714 0.1725 0.0677 51.55 71.23 72.33
Correlation 0.1517 0.1693 0.1719 0.1001
Euclidean 0.0784 0.0881 0.0874 0.0128
20 SqEuclidean 0.1109 0.1569 0.1587 0.0432 83.35 103.15 103.40
Correlation 0.1454 0.1611 0.1613 0.0793
Average percentage gain 49.86 64.93 66.33
TABLE IV: Silhouette values of different clustering algorithms for Rice.
@@ V Results
Here, we compare four clusterings. First is the standard SC as described in Section II (also natural exponential function based SC), and used in [18]. We refer to this as Old SC. Second is our proposed base “$30$” exponential function based SC as elaborated in Section III-A. We call this the Base “$30$” SC. Third is, again our proposed, base “$30$” locally scaled exponential function based SC as descibed in Section III-B. We call this New SC. Finally, the fourth is HC, which is mentioned in the literature.
@@
The results of this comparison for soybean and rice are given in Table III and IV respectively. Here, the first column denotes the number of the clusters that are chosen based upon the previous analysis. The second column contains the distance metrics used to build the similarity matrix in the clustering algorithms. Columns three through six list the Silhouette Values of the respective algorithms. Best values in a cell are highlighted in bold. Finally, columns seven through nine give the percentage gain of Old SC, Base “$30$” SC, and New SC over HC, respectively. The best values in a cell are used to compute this gain.
We conclude that the most significant improvement in clustering quality occurs when moving from Old SC to Base “30” SC with little bit more improvement when going to New SC. That is, for soybean, the gain in these three algorithms over HC is $32.15\%$, $72.74\%$, and $81.40\%$, respectively. In other words, New SC is ${\bf 35\%}$ better than Old SC.
For rice, the gain in these three algorithms over HC is $49.86\%$, $64.93\%$, and $66.33\%$, respectively. In other words, New SC is ${\bf 11\%}$ better than Old SC. To sum up, New SC yields the best results overall, with soybean showing more significant improvement than rice.
VI Conclusion and Future Work
Phenotypic data of plants is commonly used to group species into different categories, which is further used in breeding programs. Hierarchical Clustering (HC) is a common algorithm that is used for implementing such groupings. Since this algorithm is not very accurate, recently authors in [18] proposed the use of the standard Spectral Clustering (SC) to improve accuracy. They demonstrated the usefulness of their algorithm via experiments on the soybean plant.
In this work, we propose a novel base “a” locally scaled SC that improves the standard SC. First, using spectral graph theory, specifically the Cheeger’s inequality, we theoretically show that using a base “a” exponential function as the similarity function, with increasing base value, leads to improved performance of the SC algorithm. We also integrate our technique with the existing idea of “local scaling”. Second, we perform an eigenvalue analysis to support our theoretical result. Third, using extensive experiments we demonstrate usefulness of our approach. That is, on $2376$ soybean species and $1865$ rice species, our new algorithm is 35% and 11% better than the standard SC, respectively.
@@ VI Conclusion and Future Work
There are multiple future work directions here. First, in one of the seminal works [14], the authors have listed sufficiency conditions for SC to work well. It would be very useful to translate those conditions to plant data. Second, it would be useful to experiment with other accurate clusterings (e.g., see [7]). Third, although phenotypic characteristics are useful for clustering, genetic data of plant species carries more information. In our earlier work [19], we had explored the possibility of using genetic data for clustering and sampling, however, reduced data was used there. It would be interesting to experiment with the full data exhaustively.
Fourth, it would be interesting to improve the existing clustering using mathematical optimisation as in [1] and using approximate computing as in [6]. Finally, and fifth, implicit relation between phenotypic and genetic data, as done in digital libraries content here [9], could help in better clustering for both types of data.